OpenStudy (anonymous):

please some one help me 1.Finance After taxes, Alexandra’s take-home pay is 7/10 of her salary before taxes. Write and solve an equation to find Alexandra’s salary before taxes for the pay period that resulted in $392 of take-home pay. 2.Earth Science Your weight on the Moon is about 1/6 of your weight on Earth. Write and solve an equation to show how much a person weighs on Earth if he weighs 16 pounds on the Moon. How could you check that your answer is reasonable?

OpenStudy (texaschic101):

1. 7/10 of s = t , 7/10s = t where s stands for salary and t stands for take home pay. if t = 392.... (7/10)s = 392 s = 392/(7/10) s = 392 x 10/7 s = 3920/7 s = 560 Her salary was $ 560 2. m = 1/6e where m stands for moon weight and e stands for Earth weight where m = 16 16 = 1/6e 16/(1/6) = e 16 x (6/1) = e 96 = e check...16 = (1/6)96 16 = 1/6 x 96/1 16 = 96/6 16 = 16 (correct) 16 lbs. on the moon = 96 lbs. on Earth